# The ubiquity of variable radio emission and spin-down rates in pulsars

### v1

@@ 4.2 New pulsars with both profile and spin-down variations
Another pulsar that shows rapid profile shape variations in spite of a slowly changing spin-down rate. Much of the rapid changes either side of the profile peak can be attributed to epoch-to-epoch jitter, with a deficit of emission on one side being countered by an excess on the other. There is however a clear correlation between the slowed intensity changes across the profile peak itself and the spin-down rate at a lag of $\sim$1000 d. Inspection of the $\dot{\nu}$ model in Figure 4 reveals two slow, transient increases in spin-down rate that took place between MJD 53616-54441 and MJD 57564-58852 with $\delta\dot{\nu}/\dot{\nu}\sim 0.3$% and $\sim 0.2$% respectively. A small offset in the pulsar position in our timing model is likely to be responsible for the high frequency fluctuation in $\dot{\nu}$ that is superimposed on the slow variations with time.
@@
A 0.49 s radio pulsar that emits a two-component profile consisting of a bright leading peak and a trailing shoulder. Its spin-down timeseries reveals it undergoes transient decreases in spin-down rate by $\delta\dot{\nu}/\dot{\nu}\sim 0.1$-$0.9$% that last 228-975 d. Each event is separated in time by approximately 1000-2000 d. The first and fourth decreases in spin-down rate appear to coincide with an excess of emission from the trailing shoulder component, despite the overall profile variability map covering this segment being only weakly correlated with the $\dot{\nu}$ timeseries. Curiously, the recovery stage of the second spin-down event that we detect appears to also coincide with an excess in emission from the shoulder component, while the third and fifth events do not have any corresponding profile shape change.
@@
Another bright radio pulsar with profile shape changes confined to the trailing edge of its profile. Its changes appear to be weakly anti-correlated with high fluctuation frequency variations in the spin-down timeseries, for which our periodicity analysis failed to recover a corresponding (quasi-)period. Instead only the lower frequency variation with a quasi-period of $\sim$1066 d was recovered. Similar to PSR J1327--6222, our monthly observing cadence may be insufficient to fully resolve the short-time scale quasi-periodicity.
@@
A bright 22 mJy pulsar (at 1.4 GHz) with a triple component profile, where all three components display seemingly random fluctuations in relative intensity. Like many of our newly identified profile/spin-down varying pulsars, the component shape changes take place on a more rapid timescale compared to the slowly changing value of $\dot{\nu}$. Though there is a weak anti-correlation between the variability map and the spin-down rate across all three components. The Lomb-Scargle periodogram computed from the $\dot{\nu}$ timeseries recovered a quasi-period of $\sim$1015 d.
@@ 4.3 Other noteworthy pulsars
An unusual young, energetic pulsar surrounded by a wind nebula in the supernova remnant SNR G320.4--1.2 (see, Gaensler et al. 2002 and references therein). In spite of its relative youth, the long-term timing of PSR J1513--5908 is uncharacteristically stable, enabling measurements of both its braking index of $n=2.832\pm 0.003$ and higher-order braking index of $m=17.6\pm 1.9$ (Livingstone & Kaspi, 2011). It has also not glitched in over 30 years of timing (Parthasarathy et al., 2020), counter to the expected glitch rate of 0.9 yr-1 for pulsars with similar rotational properties (Lower et al., 2021). After correcting for the spin-frequency of PSR J1513--5908 and its first, second and third derivatives, our Gaussian process fit to its timing residuals reveals short timescale fluctuations in $\dot{\nu}$ with a low Q-factor quasi-period of $\sim$307 d. This short-term behaviour is superimposed on a longer timescale, smooth change in spin-down, where the amplitude of $\delta\dot{\nu}/\dot{\nu}\sim 0.03$% is the smallest out of our entire sample. No changes in profile shape were identified, owing to our relatively short integration times and its weak profile.
@@
A high magnetic field strength pulsar with an unusually low braking index of $n=1.2\pm 0.2$ (Espinoza et al., 2011; Lower et al., 2021). Subtracting the linear slope induced by the $\ddot{\nu}$ term leaves a predominantly flat spin-down timeseries for most of its 1997-2014 timing history. However, there are several transient changes in $\dot{\nu}$ occurring between MJD 53489-53729, MJD 54270-54550 and MJD 55425-55597. The second and largest event appears qualitatively similar in direction and shape to a recent transient spin-down change identified in X-ray timing of PSR J0540--6919 (B0540--69; Espinoza et al. 2024), albeit with an order of magnitude larger fractional increase in spin-down rate of $\delta\dot{\nu}/|\dot{\nu}|\sim 0.1\%$. No profile shape changes are detected alongside any of these spin-down events. However, this is not surprising given the relatively weak and scatter-broadened profile of PSR J1734--$3$ at 1.4 GHz.
@@ 5 Discussion
We have identified 238 pulsars that exhibit substantial variations in their spin-down rates, of which 52 also display changes in profile shape, meaning that we have assembled the largest sample of variable pulsars to date. The propensity of these behaviours confirms previous suggestions that radio emission and spin-down state switching is commonplace among the broader pulsar population. With this large sample on hand we can also investigate how the intensity of these effects scales with the average pulsar spin and spin-down rate, in addition to exploring the origins of specific sub-types of profile/spin-down rate variability.
@@ 5.1 Scaling relations across the population
While a direct causal relationship between the observed spin-down variability in pulsars and external or internal physical processes is yet to be achieved, we can explore how this behaviour scales phenomenologically with various pulsar properties. In Figure 11 we present the measured min/max difference in spin-down rate against the corresponding ‘weak’ (minimum) spin-down rate value for the 238 variable pulsars in our sample, alongside values from several previous works. Lyne et al. (2010) noted their sample of 17 correlated profile/spin-down state switching pulsars appeared to follow a simple power-law relation, where $\delta\dot{\nu}\propto 0.01|\dot{\nu}_{\rm weak}|$. A somewhat shallower power-law relation of $\log_{10}(\delta\dot{\nu}/|\dot{\nu}_{\rm weak}|)\propto 0.84\pm 0.02$ was recovered using extended timing of the same pulsar sample by Shaw et al. (2022). While there is a substantial number of pulsars in Figure 11 that reside about the original Lyne et al. (2010) relation (shown in black), about half of our larger sample are situated between one to three orders of magnitude below the relation. This is particularly true for pulsars with both higher and lower weak spin-down rates than those among the Lyne et al. (2010) sample.
@@
This is shallower than the aforementioned relation from Lyne et al. (2010), but consistent with that inferred by Shaw et al. (2022) at the 68 per cent confidence interval. The power-law amplitude by comparison is substantially lower, resulting from the larger number of pulsars with lower $\delta\dot{\nu}$ than those analysed by Lyne et al. (2010) and Shaw et al. (2022).
Figure 11: Relationship between spin-down variability and weak spin-down state. Values for the pulsars studied by Shaw et al. (2022) are shown in orange, the red diamond indicates PSR J0738--4042 (Lower et al., 2023), pulsars from Basu et al. (2024) are the teal hexagons, and our measurements are plotted in blue. Note, there are several pulsars in common between these datasets as indicated by overlapping points. The black line is the $\delta\dot{\nu}\propto 0.01|\dot{\nu}|_{\rm weak}$ relation from Lyne et al. (2010). The pink line is the median a-posteriori fit to the blue points with the pink shading indicating the 68 per cent and 97  per cent confidence intervals.
@@
which is close to being consistent with the same $\delta\dot{\nu}\propto|\dot{\nu}|$ scaling relation from Equation 7. The posterior distribution for the $\nu$ scaling parameter overlaps with zero at the 95 per cent confidence interval, indicating there is little dependence between the amount of spin-down variability in a given pulsar and its spin frequency.
@@ 5.2 Implications for precision timing experiments
Millisecond pulsars are known to display emission state switching on short timescales (Mahajan et al., 2018; Miles et al., 2022; Nathan et al., 2023) and transient profile shape changes (Lam et al., 2018; Goncharov et al., 2021; Jennings et al., 2024), with the latter effect requiring either profile-domain timing techniques to correct for them or careful removal of the corresponding signal in the timing residuals (see, e.g. Shannon et al. 2016). The propensity of spin-down rate variations among our pulsar sample and clear positive scaling in variability amplitude with $\dot{\nu}$ suggests this phenomena should also be both present and detectable among the millisecond pulsar population. Accurately correcting for these effects is particularly important for precision timing experiments aiming to detect the stochastic gravitational-wave background from inspiralling supermassive binary black holes.
Recent hints of such a gravitational-wave background signal among pulsar timing experiments have revealed an unexpectedly large amplitude given previous upper limits (Agazie et al., 2023; EPTA Collaboration et al., 2023; Xu et al., 2023; Reardon et al., 2023; Miles et al., 2024), which appears to be increasing with time (Reardon et al., 2023). It is presently unclear whether this is a genuine feature of the common signal, in which case it would be inconsistent with expectations for an isotropic gravitational-wave background, but could be the result of an unmodelled process intrinsic to the observed millisecond pulsars. Spin-down rate changes similar to those among our P574 pulsars would manifest as a form of non-stationary timing noise, for which the red power-law models that are typically used to account for timing noise have been demonstrated as insufficient for fully characterising the resulting timing variations (Keith & Niţu, 2023; Lower et al., 2023). While it has been claimed that non-stationary processes are not present in the European Pulsar Timing Array data (Falxa et al., 2024), the time-varying common noise signal was assumed to be accurately described by a power-law process.
Assuming the rotational characteristics of millisecond pulsars follows the same scaling as the young pulsar population, then our results can be used to predict the expected spin-down rate variability among the pulsars currently monitored by pulsar timing arrays. The spin-down rates of the pulsars included in the upcoming third data release of the International Pulsar Timing Array (K. Liu, private communication) largely range between $\dot{\nu}=-10^{-16}$ s-2 and $-10^{-14}$ s-2, alongside several noteworthy outliers such as PSR B1937$+$21 with its high spin-down rate of $-4.3\times 10^{-14}$ s-2. If the variable spin-down rate process is active among them, then from Figure 11 we would expect these pulsars to display changes in $\dot{\nu}$ with amplitudes ranging between $\delta\dot{\nu}=10^{-19}$ to $10^{-15}$ s-2. This would be easily discernable using the Gaussian process regression framework that we employed.
@@ 5.3 Quasi-periodic variability
Among our broader sample there exists a sub-group of 45 pulsar that display striking quasi-periodic oscillations in their spin-down timeseries. These pulsars were predominantly identified through a combination of visual inspection of their timing residuals, spin-down timeseries and Lomb-Scargle periodograms. This group includes those that were previously identified by Lyne et al. (2010) and Kerr et al. (2016). Modulation periods ($\mathcal{P}$) listed in Table 1 were recovered from Lomb-Scargle periodograms generated from their spin-down timeseries. We failed to recover the sub-annual periodicities of 120 d and 123 d in PSRs J1646--4346 and J1825--1446 that were found by Kerr et al. (2016). This may be due to a combination of our extended timing baselines and differences in methodology. Ten of the pulsars in Table 1 display both long and short modulation timescales that may be harmonically related. The ‘long’ modulation periods in PSRs J1136--5525, J1243--6423, J1649--4653, J1705--4306 and J1830--1059 are either close to or exactly twice the duration of the corresponding ‘short’ periods in these pulsars, and hence may be second harmonics. In comparison the dual modulation timescales in PSRs J1056--6258 J1224--6407, J1326--6700, J1600--5751 and J1722--3712 display approximate fractional relations of 5/2, 7/4, 9/5, 3/2 and 7/5 respectively.
@@
which is inconsistent with the free-precession relation. While it consistent with the $\mathcal{P}\approx 1.4\,{\rm yr}\,(P/1\,{\rm s})^{1/2}(\lambda/10^{6}\,{\rm cm})$ scaling relation predicted for Tkachenko waves when assuming an oscillation wavelength of $2\times 10^{6}$ cm at the 95 per cent confidence interval (see figure 5 in Jones 2012), this is likely just a coincidence given the scatter among our measurements.
Figure 12: Comparison of spin-down modulation period with $P$. Grey points are the dominant quasi-periods of all pulsars indicated by their Lomb-Scargle periodograms. Dark blue points indicate the highly quasi-periodic pulsars listed in Table 1, while the orange points are from Table 15 of Lyne et al. (2010). The solid black curve is the $P/\mathcal{P}=5\times 10^{-9}$ relation for freely precessing pulsars from Jones (2012), while the pink line and shading indicates our median fit along with the 68 and 95  per cent confidence intervals. The dashed horizontal line indicates a quasi-period of 1 yr.
@@
Figure 13: As Figure 12, but for pulsar characteristic age ($\tau_{c}$). The solid black line indicates the relation Equation 42 of Jones (2012) for $H_{c}=10^{15}$ G.

### v2

@@ 4.2 New pulsars with both profile and spin-down variations
The radio profile of PSR J1824--1945 displays rapid shape variations in spite of a slowly changing spin-down rate. Much of the rapid changes either side of the profile peak can be attributed to epoch-to-epoch jitter, with a deficit of emission on one side being countered by an excess on the other. There is however a clear correlation between the slowed intensity changes across the profile peak itself and the spin-down rate at a lag of $\sim$1000 d. Inspection of the $\dot{\nu}$ model in Figure 5 reveals two slow, transient increases in spin-down rate that took place between MJD 53616-54441 and MJD 57564-58852 with $\delta\dot{\nu}/|\dot{\nu}|\sim 0.3$% and $\sim 0.2$% respectively. A small offset in the pulsar position in our timing model is likely to be responsible for the high frequency fluctuation in $\dot{\nu}$ that is superimposed on the slow variations with time.
@@
The 0.49 s radio pulsar PSR J1845--0434 displays a two-component profile consisting of a bright leading peak and a trailing shoulder. Its spin-down timeseries reveals it undergoes transient decreases in spin-down rate by $\delta\dot{\nu}/|\dot{\nu}|\sim 0.1$-$0.9$% that last 228-975 d. Each event is separated in time by approximately 1000-2000 d. The first and fourth decreases in spin-down rate appear to coincide with an excess of emission from the trailing shoulder component, despite the overall profile variability map covering this segment being only weakly correlated with the $\dot{\nu}$ timeseries. Curiously, the recovery stage of the second spin-down event that we detect appears to also coincide with an excess in emission from the shoulder component, while the third and fifth events do not have any corresponding profile shape change.
@@
PSR J1913--0440 is another bright radio pulsar with profile shape changes confined to the trailing edge of its profile. Its changes appear to be weakly anti-correlated with high fluctuation frequency variations in the spin-down timeseries, for which our periodicity analysis failed to recover a corresponding (quasi-)period. Instead only the lower frequency variation with a quasi-period of $\sim$1066 d was recovered. Similar to PSR J1327--6222, our monthly observing cadence may be insufficient to fully resolve the short-time scale quasi-periodicity.
@@
This bright 22 mJy pulsar (at 1.4 GHz) with a triple component profile, where all three components display seemingly random fluctuations in relative intensity. Like many of our newly identified profile/spin-down varying pulsars, the component shape changes take place on a more rapid timescale compared to the slowly changing value of $\dot{\nu}$. However, there is a weak anti-correlation between the variability map and the spin-down rate across all three components. The Lomb-Scargle periodogram computed from the $\dot{\nu}$ timeseries recovered a quasi-period of $\sim$1015 d.
@@ 4.3 Other noteworthy pulsars
An unusual young, energetic pulsar surrounded by a wind nebula in the supernova remnant SNR G320.4--1.2 (see, Gaensler et al. 2002 and references therein). In spite of its relative youth, the long-term timing of PSR J1513--5908 is uncharacteristically stable, enabling measurements of both its braking index of $n=2.832\pm 0.003$ and higher-order braking index of $m=17.6\pm 1.9$ (Livingstone & Kaspi, 2011). It has also not glitched in over 30 years of timing (Parthasarathy et al., 2020), counter to the expected glitch rate of 0.9 yr-1 for pulsars with similar rotational properties (Lower et al., 2021). After correcting for the spin-frequency of PSR J1513--5908 and its first, second and third derivatives, our Gaussian process fit to its timing residuals reveals short timescale fluctuations in $\dot{\nu}$ with a low Q-factor quasi-period of $\sim$307 d. This short-term behaviour is superimposed on a longer timescale, smooth change in spin-down, where the amplitude of $\delta\dot{\nu}/|\dot{\nu}|\sim 0.03$% is the smallest out of our entire sample. No changes in profile shape were identified, owing to our relatively short integration times and its weak profile.
@@
This high magnetic field strength pulsar with an unusually low braking index of $n=1.2\pm 0.2$ (Espinoza et al., 2011; Lower et al., 2021). Subtracting the linear slope induced by the $\ddot{\nu}$ term leaves a predominantly flat spin-down timeseries for most of its 1997-2014 timing history. However, there are several transient changes in $\dot{\nu}$ occurring between MJD 53489-53729, MJD 54270-54550 and MJD 55425-55597. The second and largest event appears qualitatively similar in direction and shape to a recent transient spin-down change identified in X-ray timing of PSR J0540--6919 (B0540--69; Espinoza et al. 2024), albeit with an order of magnitude larger fractional increase in spin-down rate of $\delta\dot{\nu}/|\dot{\nu}|\sim 0.1\%$. No profile shape changes are detected alongside any of these spin-down events. However, this is not surprising given the relatively weak and scatter-broadened profile of PSR J1734--$3$ at 1.4 GHz.
@@ 5 Discussion
We have identified 238 pulsars that exhibit substantial variations in their spin-down rates, of which 52 also display changes in profile shape, meaning that we have assembled the largest sample of variable pulsars to date. The propensity of these behaviours confirms previous suggestions that radio emission and spin-down state switching is commonplace among the broader pulsar population. With this large sample on hand we can also investigate how the amplitude of these effects scales with the average pulsar spin and spin-down rate, in addition to exploring the origins of specific sub-types of profile/spin-down rate variability.
@@ 5.1 Scaling relations across the population
While a direct causal relationship between the observed spin-down variability in pulsars and external or internal physical processes is yet to be established, we can explore how this behaviour scales phenomenologically with various pulsar properties. In Figure 12 we present the measured min/max difference in spin-down rate against the corresponding ‘weak’ (minimum) spin-down rate value for the 238 variable pulsars in our sample, alongside values from several previous works. Lyne et al. (2010) noted their sample of 17 correlated profile/spin-down state switching pulsars appeared to follow a simple power-law relation, where $\delta\dot{\nu}\propto 0.01|\dot{\nu}_{\rm weak}|$. A somewhat shallower power-law relation of $\log_{10}(\delta\dot{\nu}/|\dot{\nu}_{\rm weak}|)\propto 0.84\pm 0.02$ was recovered using extended timing of the same pulsar sample by Shaw et al. (2022). While there is a substantial number of pulsars in Figure 12 that reside about the original Lyne et al. (2010) relation (shown in black), about half of our larger sample are situated between one to three orders of magnitude below the relation. This is particularly true for pulsars with both higher and lower weak spin-down rates than those among the Lyne et al. (2010) sample.
@@
This is shallower than the aforementioned relation from Lyne et al. (2010), but consistent with that inferred by Shaw et al. (2022) at the 68% confidence interval. The power-law amplitude by comparison is substantially lower, resulting from the larger number of pulsars with lower $\delta\dot{\nu}$ than those analysed by Lyne et al. (2010) and Shaw et al. (2022).
Figure 12: Relationship between spin-down variability and weak spin-down state. Values for the pulsars studied by Shaw et al. (2022) are shown in orange, the red diamond indicates PSR J0738--4042 (Lower et al., 2023), pulsars from Basu et al. (2024) are the teal hexagons, and our measurements are plotted in blue. Note, there are several pulsars in common between these datasets as indicated by overlapping points. The black line is the $\delta\dot{\nu}\propto 0.01|\dot{\nu}|_{\rm weak}$ relation from Lyne et al. (2010). The pink line is the median a-posteriori fit to the blue points with the pink shading indicating the 68% and 97% confidence intervals.
@@
which is close to being consistent with the same $\delta\dot{\nu}\propto|\dot{\nu}|$ scaling relation from Equation 7. The posterior distribution for the $\nu$ scaling parameter overlaps with zero at the 95% confidence interval, indicating there is little dependence between the amount of spin-down variability in a given pulsar and its spin frequency.
@@ 5.2 Implications for precision timing experiments
Millisecond pulsars are known to display emission state switching on short timescales (Mahajan et al., 2018; Miles et al., 2022; Nathan et al., 2023) and transient profile shape changes (Shannon et al., 2016; Lam et al., 2018; Goncharov et al., 2021; Jennings et al., 2024), with the latter effect requiring either profile-domain timing techniques to correct for them or careful removal of the corresponding signal in the timing residuals (see, e.g. Shannon et al. 2016). The propensity of spin-down rate variations among our pulsar sample and clear positive scaling in variability amplitude with $\dot{\nu}$ suggests this phenomena should also be both present and detectable among the millisecond pulsar population. Accurately correcting for these effects is particularly important for precision timing experiments aiming to detect the stochastic gravitational-wave background.
Recent hints of such a gravitational-wave background signal among pulsar timing experiments have revealed an unexpectedly large amplitude given previous upper limits (Agazie et al., 2023a; EPTA Collaboration et al., 2023b; Xu et al., 2023; Reardon et al., 2023; Miles et al., 2025b), which appears to be increasing with time (Reardon et al., 2023). It is presently unclear whether this is a genuine feature of the common signal, in which case it would be inconsistent with expectations for an isotropic gravitational-wave background, but could be the result of an unmodelled process intrinsic to the observed millisecond pulsars. Spin-down rate changes similar to those among our P574 pulsars would manifest as a form of non-stationary timing noise, for which the red power-law models that are typically used to account for timing noise have been demonstrated as insufficient for fully characterising the resulting timing variations (Keith & Niţu, 2023; Lower et al., 2023).
Assuming the rotational characteristics of millisecond pulsars follows the same scaling as the young pulsar population, then our results can be used to predict the expected spin-down rate variability among the pulsars currently monitored by pulsar timing arrays. The spin-down rates of the pulsars included in the upcoming third data release of the International Pulsar Timing Array (K. Liu, private communication; Zic et al. 2023; EPTA Collaboration et al. 2023a; Agazie et al. 2023b; Miles et al. 2025a) largely range between $\dot{\nu}=-10^{-16}$ s-2 and $-10^{-14}$ s-2, alongside several noteworthy outliers such as PSR B1937$+$21 with its high spin-down rate of $-4.3\times 10^{-14}$ s-2. If the variable spin-down rate process is active among them, then from Figure 12 we would expect these pulsars to display changes in $\dot{\nu}$ with amplitudes ranging between $\delta\dot{\nu}=10^{-19}$ to $10^{-15}$ s-2. This would be easily discernable using the Gaussian process regression framework that we employed.
@@ 5.3 Quasi-periodic variability
Among our broader sample there exists a sub-group of 45 pulsars that display striking quasi-periodic oscillations in their spin-down timeseries. These pulsars were predominantly identified through a combination of visual inspection of their timing residuals, spin-down timeseries and Lomb-Scargle periodograms. This group includes five that were previously identified by Lyne et al. (2010) and Kerr et al. (2016). Modulation periods ($\mathcal{P}$) listed in Table 1 were recovered from Lomb-Scargle periodograms generated from their spin-down timeseries. We failed to recover the sub-annual periodicities of 120 d and 123 d in PSRs J1646--4346 and J1825--1446 that were found by Kerr et al. (2016). This may be due to a combination of our extended timing baselines and differences in methodology. Ten of the pulsars in Table 1 display both long and short modulation timescales that may be harmonically related. The ‘long’ modulation periods in PSRs J1136--5525, J1243--6423, J1649--4653, J1705--4306 and J1830--1059 are either close to or exactly twice the duration of the corresponding ‘short’ periods in these pulsars, and hence may be second harmonics. In comparison the dual modulation timescales in PSRs J1056--6258 J1224--6407, J1326--6700, J1600--5751 and J1722--3712 display approximate fractional relations of 5/2, 7/4, 9/5, 3/2 and 7/5 respectively.
@@
which is inconsistent with the free-precession relation. While it is consistent with the $\mathcal{P}\approx 1.4\,{\rm yr}\,(P/1\,{\rm s})^{1/2}(\lambda/10^{6}\,{\rm cm})$ scaling relation predicted for Tkachenko waves when assuming an oscillation wavelength of $2\times 10^{6}$ cm at the 95% confidence interval (see figure 5 in Jones 2012), this is likely just a coincidence given the scatter among our measurements.
Figure 13: Comparison of spin-down modulation period with $P$. Grey points are the dominant quasi-periods of all pulsars indicated by their Lomb-Scargle periodograms. Dark blue points indicate the highly quasi-periodic pulsars listed in Table 1, while the orange points are from Table 15 of Lyne et al. (2010). The solid black curve is the $P/\mathcal{P}=5\times 10^{-9}$ relation for freely precessing pulsars from Jones (2012), while the pink line and shading indicates our median fit along with the 68% and 95% confidence intervals. The dashed horizontal line indicates a quasi-period of 1 yr.
@@
Figure 14: As Figure 13, but for pulsar characteristic age ($\tau_{c}$). The solid black line indicates the relation from equation 42 of Jones (2012) for $H_{c}=10^{15}$ G.